# Make the Fastest Faster: Importance Mask for Interactive Volume Visualization using Reconstruction Neural Networks

### v1

@@ 4 Limitations and Future Work
One limitation of the current design is that the mean value used to determine the percentage of C-Image as important pixels is a predefined hyperparameter. While users can adjust the mean value during inference, our experiments indicate that the network trained with a specific mean value $m$ achieves the best reconstruction quality when the inference mean value is also close to $m$. Using a mean value significantly higher than $m$, resulting in a large number of pixels being rendered as important pixels, does not improve reconstruction quality and may even lead to worse results. It would be highly beneficial to develop a method that enables the network to also learn an adaptive mean value, resulting in a dynamic percentage of important pixels in the C-Image that adjusts based on the underlying rendering results. This approach can also assist in quickly identifying the optimal ORP for a given reconstruction quality error tolerance.
Various aspects of the pipeline need detailed investigation in the future. In this work, we only investigate the static fixed foveated rendering, it is worth investigating whether dynamic foveal rendering could impact reconstruction quality, rendering latency, and the pattern of the IM learned. IML Net is trained only using pixel-wise loss (MSE) in our work, other loss configurations, like structural similarity-based loss (SSIM) and perceptual loss, can be introduced to evaluate how the loss function affects the pattern of the learned IM. It is worth investigating which factors influence the distribution of important pixels and how many pixels are needed to render a specific frame. Factors such as image complexity, region size, and color dynamics may all contribute to determining which pixels are selected as important pixels. We also see the potential of utilizing the proposed IML Net to generate IMs with various levels of detail (LoD), enabling the construction of a multi-resolution representation to further optimize rendering performance.
@@ 5 Conclusion
In this work, we present IML Net and IMS Net to learn and generate an important mask that will help to further improve the already fast volume visualization using a reconstruction neural network. Our work represents the first attempt to optimize predefined sampling patterns used in various reconstruction neural networks through a unified interface, utilizing compaction and decompaction layers. We showcase the improvement of rendering latency using our method on two state-of-the-art reconstruction neural networks, EnhanceNet and FoVolNet. Using four distinct volumetric datasets, we quantitatively measured how much rendering latency can be reduced for free without compromising noticeable reconstruction quality. Our proposed IML Net is also flexible to be trained independently from the downstream reconstruction neural network, allowing quick and practical application of our method on top of an off-the-shelf pre-trained reconstruction neural network.

### v2

@@ 4 Limitations and Future Work
One limitation of the current design is that the mean value used to determine the percentage of C-Image as important pixels is a predefined hyperparameter. While users can adjust the mean value during inference, our experiments indicate that the network trained with a specific mean value $m$ achieves the best reconstruction quality when the inference mean value is also close to $m$. Using a mean value significantly higher than $m$, resulting in a large number of pixels being rendered as important pixels, does not improve reconstruction quality and may even lead to worse results. It would be highly beneficial to develop a method that enables the network to also learn an adaptive mean value, resulting in a dynamic percentage of important pixels in the C-Image that adjusts based on the underlying rendering results. This approach can also assist in quickly identifying the ORP for a given reconstruction quality error tolerance. More comprehensive methods of identifying ORP, rather than setting a 1 dB threshold, are also worth exploring.
Various aspects of the pipeline need detailed investigation in the future. In this work, we only investigate the static fixed foveated rendering, it is worth investigating whether dynamic foveal rendering could impact reconstruction quality, rendering latency, and the pattern of the IM learned. IML Net is trained only using pixel-wise loss (MSE) in our work, other loss configurations, like structural similarity-based loss (SSIM), can be introduced to evaluate how the loss function affects the pattern of the learned IM. It is worth investigating which factors influence the distribution of important pixels and how many pixels are needed to render a specific frame. Factors such as image complexity, region size, and color dynamics may all contribute to determining which pixels are selected as important pixels. We also see the potential of utilizing the proposed IML Net to generate IMs with various levels of detail (LOD), enabling the construction of a multi-resolution representation to further optimize rendering performance.
@@ 5 Conclusion
In this work, we present IML Net and IMS Net to learn and synthesize an important mask that will help to further improve the already fast volume visualization using a reconstruction neural network. Our work presents the first attempt to directly synthesize an importance mask from a given view for predefined sampling patterns used in various reconstruction neural networks, via a unified interface that leverages the proposed compaction and decompaction layers. We showcase the quantitative improvement of rendering latency using our method on two state-of-the-art reconstruction neural networks, EnhanceNet and FoVolNet, on four distinct volumetric datasets, without noticeable degradation of reconstruction quality. Our proposed IML Net is also flexible to be trained independently from the downstream reconstruction neural network, allowing quick and practical application of our method on top of various off-the-shelf pre-trained reconstruction neural networks.
Acknowledgement
This research has been sponsored in part by the National Science Foundation grants IIS-1423487 and IIS-1652846, and Advanced Scientific Computing Research, Office of Science, U.S. Department of Energy, under contracts DE-AC02-06CH11357, program manager Hal Finkel. The authors would like to thank the anonymous reviewers for their insightful comments.